Tiffany claims she is older than Nicole by only 2 years.
Nicole claims she is half of Jennifer's age.
Jennifer admits she is the oldest but is older than Tiffany by only 17 years.
What is each womans age?

Answer by AnlytcPhil(1278)

Tiffany claims she is older than Nicole by only 2 years.
Nicole claims she is half of Jennifer's age.
Jennifer admits she is the oldest but is older than Tiffany
by only 17 years.
What is each womans age?

>>...Tiffany...is older than Nicole by...2 years...<<

 T = N + 2


>>...Nicole...is half of Jennifer's age...<<

 N = 1/2J

>>...Jennifer...is older than Tiffany by...17 years...<<

 J = T + 17

So we have a system three equations in 3 unknowns:

T = N + 2
N = 1/2J
J = T + 17

Can you solve that system?  If not post again:

Answer = J = 38, N = 19, T = 21

Edwin McCravy
AnlytcPhil@aol.com